There are compressed isothermally 22 m3 of air measured at 27oC and 1378 kPa. If p2 = 3790 kPa, compute (a) the heat (gain or loss?), (b) the change in entropy.


1
Expert's answer
2021-12-06T09:40:37-0500

"V_2=V_1\frac{p_1}{p_2}=8~m^3,"

a)

"Q=A=p_1V_1\ln\frac{V_2}{V_1}=31~MJ,"

b)

"\Delta S=\frac Q{T_1}=102~\frac{kJ}K."